We know this equation A°=1.Prove this?

Grade:12

3 Answers

Prangan Bhattacharyya
11 years ago

we know that    A^m/A^m=1

                   so A^(m-m)=1

                   so A^0=1.....hope this helps

Akul Goel
11 years ago

we know that log1 to any base =0 take antilog threfore a0 =1

Raghuvaran varan Chandragiri
11 years ago

Mathematics Assignment Help, We know this equation a°=1.prove this?, we know that A^m/A^m=1 so A^(m-m)=1 so A^0=1